What are the odds of getting a five and a jack out of a deck of 52 cards?
I thought it was 8/52 but that was wrong.

Found 3 solutions by Alan3354, scott8148, oberobic:
Answer by Alan3354(69443)

What are the odds of getting a five and a jack out of a deck of 52 cards?
-----------
It's 2 draws from a deck of 52, it's
4/52 * 4/52 = 1/169 (called with replacement)
----------
If it's drawing a card, keeping it, and drawing another, it's
4/52 * 4/51 = 4/663 (without replacement)

Answer by scott8148(6628)

there are 16 ways (4 * 4) to get a five and a jack ___ any of 4 fives with any of 4 jacks

there are 52C2 ways of getting two cards


HINT: probability and odds are NOT exactly the same

Answer by oberobic(2304)

With 52 cards, the chance of getting a five is 4/42 = 1/13.
With 52 cards, the chance of getting a jack is 4/52 = 1/13.
.
When you are asked an "and" question regarding probabilities, it means "multiply."
.
The chance of getting a 5 and then a jack on successive draws, with replacement is 1/13*1/13 = 1/169.
.
If you do not replace the five, then your chance of drawing the five is 4/52=1/13 and drawing a jack after drawing a five is 4/51.
1/13*4/51 = 4/663